b3.4 2018 Individual Tax Rates Single Individuals If Your Taxable Income Is You Pay This Amount...

b3.4

2018 Individual Tax Rates
Single Individuals

If Your Taxable Income Is
You Pay This
Amount on the
Base of the Bracket
Plus This Percentage
on the Excess over the
Base (Marginal Rate)
Average Tax
Rate at
Top of Bracket
Up to $9,525 $0 10.0 % 10.0 %
$9,525 - $38,700 952.50 12.0 11.5
$38,700 - $82,500 4,453.50 22.0 17.1
$82,500 - $157,500 14,089.50 24.0 20.4
$157,500 - $200,000 32,089.50 32.0 22.8
$200,000 - $500,000 45,689.50 35.0 30.1
Over $500,000 150,689.50 37.0 37.0

Standard deduction for individual: $12,000

Married Couples Filing Joint Returns

If Your Taxable Income Is
You Pay This
Amount on the
Base of the Bracket
Plus This Percentage
on the Excess over the
Base (Marginal Rate)
Average Tax
Rate at
Top of Bracket
Up to $19,050 $0 10.0 % 10.0 %
$19,050 - $77,400 1,905.00 12.0 11.5
$77,400 - $165,000 8,907.00 22.0 17.1
$165,000 - $315,000 28,179.00 24.0 20.4
$315,000 - $400,000 64,179.00 32.0 22.8
$400,000 - $600,000 91,379.00 35.0 26.9
Over $600,000 161,379.00 37.0 37.0

Standard deduction for married couples filing jointly: $24,000

Quantitative Problem: Jenna is a single taxpayer. During 2018, she earned wages of $120,000. She doesn't itemize deductions, so she will take the standard deduction to calculate 2018 taxable income. In addition, during the year she sold common stock that she had owned for five years for a net profit of $7,200. How much does Jenna owe to the IRS for taxes? Do not round intermediate calculations. Round your answer to the nearest cent.

$  
Answer #1

Answer:

Tax Year: 2018

Filing status: Single

She sold common stock she had owned for 5 years for net profit of $7,200.

Hence:

Her long term capital gain = $7,200

As such:

Her gross income = 120000 + 7200 = $127,200

Her AGI = 127200 - 0 = $127,200

Her taxable income = AGI - Standard deduction = 127200 - 12000 = $115,200

Long term capital gain = $7,200

Ordinary Income = 115200 - 7200 = $108,000

Long term capital gain tax rate for her will be = 15%

Hence:

Her tax liability =14089.50 + (108000 - 82500) * 24% + 7200 * 15%

= $21,289.50

Amount Jenna owe to the IRS for taxes = $21,289.50